a square window has an of 96 square inches. what is the length of each side of the window? be sure your answer is in simplest radical form. show as much work as possible.
*can you also show me how to check the answer to make sure it is correct?
thanks for all your help!

Answer by rothauserc(4718)

The area(A) of a square is equal to the length of a side(s) squared, therefore
96 = s^2
s = square root of 96
s = 9.797
use hundredths to keep from going over 96
s = 9.79 inches
check answer
96 = 9.79^2 = 95.844 approx 96